vdoggsavi7:

Convert 5.764764764 ... to a rational expression in the form of a over b , where b ≠ 0. (4 points) a 5 and 764 over 999 b 5764 over 9999 c 5 and 764 over 99 d 5 and 999 over 764

Vocaloid:

5.764764764 let's let x = 0.764764764... (the repeating part only) therefore, 1000x = 764.764764764... notice what happens when we subtract 1000x - x = 999x = 764 (the repeating parts from both cancel out) if 999x = 764, then we can divide both sides by 999 to get x = 764/999 so we have now converted the repeating decimal to a fraction. going back to the original number, it's 5.764764764 = 5 + 764/999